OpenStudy (anonymous):

Decide whether or not the ordered pair is a solution of the equation. 2x - 4y = 22 (3,4) A. Yes B. No

OpenStudy (anonymous):

B.NO

OpenStudy (unklerhaukus):

if \((3,4)\) was a solution \[2x - 4y = 22\] \(2\times(3)-4\times(4) \) would equal \(=22\)

OpenStudy (anonymous):

maybe.... maybe not...

OpenStudy (anonymous):

the left hand side comes to -10 n the right hand side 22. so not a solution

OpenStudy (anonymous):

I take back my answer!

OpenStudy (anonymous):

I think it no too :D

OpenStudy (anonymous):

thanks for your input :D

OpenStudy (anonymous):

Challenge every answer...never know when someone will make a mistake..